# Teaching Large Language Models an Unseen Language on the Fly

### v1

@@ Abstract
Existing large language models struggle to support numerous low-resource languages, particularly the extremely low-resource ones where there is minimal training data available for effective parameter updating. We thus investigate whether LLMs can learn a new language on the fly solely through prompting. To study this question, we collect a research suite for Zhuang, a language supported by no LLMs currently. We introduce DiPMT++, a framework for adapting LLMs to unseen languages by in-context learning. Using a dictionary and only 5K parallel sentences, DiPMT++ significantly enhances the performance of GPT-4 from 0 to 16 BLEU for Chinese-to-Zhuang translation and achieves 32 BLEU for Zhuang-to-Chinese translation. Furthermore, we demonstrate the practical utility of this framework in aiding humans to translate completely unseen languages, which could contribute to the preservation of linguistic diversity.
@@ 1 Introduction
Existing large language models (LLMs) provide robust support for many high-resource languages, but their support for numerous low-resource languages is limited Ahuja et al. (2023). To adapt LLMs to low-resource languages, continual pre-training or adaptors are commonly employed Pfeiffer et al. (2020); Yong et al. (2023). However, a corpus of only a few thousand sentences is insufficient for extremely low-resource languages to update the model parameters effectively Joshi et al. (2020). Considering the inductive and mimicking capabilities of current LLMs, an interesting research question arises: Can LLMs learn a new low-resource language on the fly solely through prompting? This learning paradigm could enable more efficient utilization of limited resources and holds significant potential in areas such as language preservation and education.
@@
To explore this question, we choose Zhuang (ISO 639-1: za; ISO 639-3: zha), an extremely low-resource language, as our focus. It is the language spoken by the Zhuang people of Southern China. [^1] See more information about Zhuang in Appendix A. There are no open-source natural language processing (NLP) datasets in Zhuang, and existing LLMs do not support this language. Therefore, we curate ZhuangBench, a research suite for Zhuang, comprising a dictionary, a parallel corpus, and a machine translation test set. ZhuangBench is not only a valuable linguistic resource for this extremely low-resource language but also a challenging benchmark for LLMs, with which we can investigate how models learn an entirely new language via prompts.
We mainly focus on the machine translation (MT) task in ZhuangBench. Previous research such as DiPMT Ghazvininejad et al. (2023) has explored translation in low-resource languages via prompting, providing translations for rare words with existing dictionaries. This method is tailored for languages where LLMs already possess fundamental capabilities, especially basic syntactic knowledge, which might be learned in the form of $n$-gram language modeling. However, when confronted with an entirely new language where LLMs need to acquire its vocabulary and grammar from scratch, previous prompting methods hardly work.
In this work, we introduce DiPMT++, a framework to efficiently adapt LLMs to an unseen language via in-context learning (ICL). Built upon DiPMT, we provide models with the meanings of words appearing in the source sentence, as illustrated in Figure 1. Given the inherent incompleteness of the dictionary for low-resource languages, we enhance the lexical coverage by revisiting traditional techniques like bilingual lexicon induction and synonym expansion. To aid models in grasping basic syntax, we retrieve closely related exemplars from parallel corpora to construct the ICL prompt.
@@
To investigate the applicability of DiPMT++ in realistic scenarios, we conduct a user study of unseen language translation. We recruit participants who have no knowledge of Zhuang, and ask them to conduct translation with the given linguistic sources. Experiments reveal that if we assist humans with DiPMT++, their translation improves in both quality and efficiency. This shows our framework’s great potential in preserving underrepresented languages.
Our contributions are as follows: (1) We present ZhuangBench, a challenging benchmark for LLMs to translate an unseen language with limited linguistic resources. (2) We develop an ICL framework DiPMT++ for on-the-fly language learning, which has proven effective on two benchmarks, and explore more strategies for enhancing LLMs’ acquisition of lexical and syntactic knowledge. (3) We showcase that DiPMT++ can assist humans in translating unseen languages, which could benefit the preservation of linguistic diversity. Our code and data are available to the public [^2] https://github.com/luciusssss/ZhuangBench.
@@ 2 Related Works
Adapting LLMs to Low-Resource Languages. Continual pretraining on monolingual texts is a common practice to adapt LLMs to low-resource languages Yong et al. (2023). Techniques such as MAD-X Pfeiffer et al. (2020, 2021) and LoRA Hu et al. (2021) are used to improve the training efficiency. Yang et al. (2023) design a high- to low-resource curriculum for multilingual continual pertaining. Purkayastha et al. (2023) attempt to romanizing the unseen scripts for language adaptation.
Another line of work directly adapts instruction-tuned LLMs through supervised fine-tuning (SFT) by constructing cross-lingual instructions Cahyawijaya et al. (2023) or leveraging pivot languages Zhang et al. (2023b). Yong et al. (2023) find that SFT is sometimes more efficient for language adaptation than continual pretraining.
Unlike previous works, we explore the possibility of on-the-fly language adaptation through prompting, given limited linguistic resources.
Machine Translation with LLMs. Recent works demonstrate the effectiveness of prompting an LLM with a few MT examples Zhang et al. (2023a); Vilar et al. (2023); Garcia et al. (2023). Following them, researchers explore different strategies for ICL exemplar selection, from the perspectives of term recall Agrawal et al. (2023), knowledge relevance He et al. (2023), and cultural awareness Yao et al. (2023). However, these strategies are primarily tailored for high-resource languages.
Robinson et al. (2023) show that LLMs still fail to translate low-resource languages properly. Existing methods designed for low-resource languages include incorporating dictionaries Ghazvininejad et al. (2023); Elsner and Needle (2023), using high-resource languages as pivots Jiao et al. (2023), and refining results from smaller MT systems Cheng et al. (2023). In contrast to them, our work tackles an even more challenging scenario: translating languages that are entirely unseen by LLMs.
@@ 3 Dataset: ZhuangBench
Dictionary. The Zhuang-Chinese dictionary is collected from an online dictionary site [^3] https://zha_zho.en-academic.com/, with 16,031 Zhuang words. The average number of senses for each word is 1.4. We also convert it to a Chinese-Zhuang dictionary with 13,618 Chinese words, with an average of 2.2 translations per word.
Parallel Corpus. The parallel corpus contains 4,944 Zhuang-Chinese sentence pairs from multiple sources. 2,135 pairs are obtained from the Chinese and Zhuang versions of the Government Work Reports in China, where we map each Chinese sentence to its corresponding Zhuang translation. 2,127 pairs are collected from Zhuang textbooks. The remaining 682 pairs are example sentences in the dictionary. See more statistics in Appendix B.
Translation Test Set. We provide a hold-out test set with 200 sentence pairs for evaluation. It consists of three difficulty levels: 75 easy instances, 60 medium instances, and 65 hard instances.
The easy subsets are composed of sentences from an elementary textbook. Other subsets are collected from the official Zhuang Language Proficiency Test (Vahcuengh Sawcuengh Suijbingz Gaujsi, V.S.S.G.) in China, which is divided into three levels: elementary, intermediate, and advanced. The translation questions from the elementary and intermediate levels of V.S.S.G. form our medium subset while those from the advanced level constitute our hard subset. In Appendix B we list statistics and examples for each level. The three subsets of different difficulties exhibit a gradual increase in vocabulary coverage and sentence complexity.
@@ 4 Method
We introduce DiPMT++, a language-agnostic framework to efficiently adapt LLMs to an unseen language. It can serve as a strong baseline for the machine translation task in ZhuangBench.
@@ 4.1 Preliminary: DiPMT
DiPMT Ghazvininejad et al. (2023) is a prompting-based method for low-resource language translation. Given a source sentence, the model looks for dictionary entries for rare words in it, and adds them to the prompt directly with the format in this context, the word “[source word]” means “[target word]”. The prompt is adopted in an in-context learning manner, with $k$ demonstrations before the current testing instance.
DiPMT is designed for languages that current models perform moderately well (10 - 30 BLEU points). The authors claim that DiPMT is not suitable for languages where current models perform poorly (< 10 BLEU points), as they assume that the performance for those ones is too low to expect reasonable translations even when incorporating external information.
@@ 4.2 Our Method: DiPMT++
One may ask what we can do to actually help LLMs understand those extremely low-resource languages, for which we only have a parallel corpus of a few thousand sentences. Contrary to the pessimistic view of Ghazvininejad et al. (2023), we hypothesize that it is possible to teach LLMs a new language solely through prompting, considering LLMs’ impressive ability to infer and mimic. We make extensions to the DiPMT framework so that it can be applied to extremely low-resource languages. Leveraging the powerful reasoning capabilities of LLMs, we propose DiPMT++, a new method that allows LLMs to understand a completely new language with minimal resources.
Following DiPMT, we cast the on-the-fly machine translation as an ICL task and incorporate knowledge from bilingual dictionaries. DiPMT++ makes two key modifications to DiPMT . See the prompt template in Appendix C.
Improved Lexical Coverage. DiPMT only provides meanings for less frequent words in the sentence. For an unseen language, we need to provide translations for as many words in the sentence as possible to the model. However, not all words can be found in the dictionary for a low-resource language. DiPMT++ attempts to improve lexical coverage of the prompt by revisiting traditional statistical methods and linguistic resources. Specifically, we use the following three strategies:
@@
$\bullet$ Bilingual Lexicon Induction: Even a small-scale parallel corpus might contain words that are not included in the dictionary. Traditional statistical methods such as GIZA++ Och and Ney (2003) can efficiently mine bilingual lexicon from the corpus, which could complement the dictionary.
@@
Syntactically-Informed Exemplar. In DiPMT, the exemplars are fixed and have limited relevance to the testing instance. Their function is only to demonstrate the task. It hardly works when the LLMs have little knowledge about the grammar of an unseen language. DiPMT++ attempts to dynamically select exemplars with higher relevance and encourage models to infer elementary syntactic information from the exemplars. For a testing instance, its exemplars are retrieved from a parallel corpus. Although there are advanced retrievers such as DPR Karpukhin et al. (2020), they require training and do not support extremely low-resource languages like Zhuang. Therefore, we apply BM25 Robertson et al. (2009), a language-agnostic retrieval algorithm.
Model Chinese normal-$\to$ Zhuang (BLEU / chrF) Zhuang normal-$\to$ Chinese (BLEU / chrF)
@@
Table 1: Performance of different methods on the test set of ZhuangBench. We use 3-shot exemplars for LLM-based methods. The best scores are made bold, with the second underlined.
@@ 5.1 Experimental Setup
Backbone Models. We use three types of models as the backbone of DiPMT++: (1) LLaMA-2-chat Touvron et al. (2023), an open-source English-centric model, (2) Qwen-chat Bai et al. (2023), a bilingual model for English and Chinese, and (3) GPT-3.5-Turbo and GPT-4 OpenAI (2023), two commercial multilingual models [^4] We use the latest versions before paper submission, i.e., gpt-3.5-turbo-0125 and gpt-4-0125-preview..
Baselines. We adopt a variety of baselines for comparison. (1) Finetune: Finetuning models with parallel sentences. (2) Direct: Directly asking LLMs to perform translations without providing ICL exemplars, which, to some extent, reflects whether the LLM already knows the language. (3) DiPMT Ghazvininejad et al. (2023): Using the original design of DiPMT for LLM prompting.
Metrics. We use BLEU Papineni et al. (2002) and chrF Popović (2015), implemented by Post (2018).. BLEU is a word-level metric while chrF focuses on the character level.
Tokenization. All the models used in our experiments support Chinese. Because Zhuang adopts a Latin script, these models can tokenize Zhuang texts into subwords, or characters at least, without producing UNK. For example, LLaMA-2’s tokenizer tokenizes Liuz coengmingz. into [‘_Li’, ‘uz’, ‘_co’, ‘eng’, ‘ming’, ‘z’, ‘.’].
@@ 5.2 Results and Analyses
In Table 1, we report the results on the Chinese-to-Zhuang (zh2za) and Zhuang-to-Chinese (za2zh) translation task of ZhuangBench. See samples of outputs from different models in Appendix D.
@@
Finetuning vs. Prompting. Finetuning pretrained models is a common practice for low-resource machine translations Adelani et al. (2022). Fintuned smaller models like mT5 still have BLEU scores close to zero for Zhuang [^5] Although mT5-large achieves 15.1 chrF on zh2za, the model outputs are almost non-sense, as shown by the example in Appendix D. As Zhuang uses a Latin script with 26 characters, even a meaningless sequence would likely have a non-zero chrF score.. Through finetuning, LLaMA-2-7B-chat can develop a basic understanding of Zhuang. It performs particularly well on the easy subset of zh2za translation but still struggles with more challenging instances. We refrain from finetuning larger models due to the substantial computational resources required.
@@
Compared to the high expense of finetuning, prompting with DiPMT++ requires no training while delivering comparable or even superior performance with larger models. Prompting LLaMA-2-7B-chat with DiPMT++ only lags behind finetuning by 0.6 BLEU on zh2za and by 1.6 BLEU on za2zh. Despite having little prior knowledge about Zhuang, as evidenced by their poor performance in direct prompting, GPT-3.5 and GPT-4 achieve excellent results on translation tasks of varying difficulty levels with the assistance of DiPMT++. For instance, GPT-4 paired with DiPMT++ achieves a remarkable BLEU score of 31.9 on za2zh, which might be qualified for practical use.
@@
DiPMT vs. DiPMT++. We compare DiPMT and DiPMT++ with Qwen-chat as the backbone model. The original DiPMT, although useful for mid-source languages, has limited ability to assist LLMs in understanding a completely new language. Even Qwen-72B-chat, the largest version of Qwen-chat achieves only 5.1 BLEU on Chinese-to-Zhuang translation. After introducing two simple extensions, DiPMT++ activate the reasoning ability of LLMs and greatly boost the performance. For example, DiPMT++ increases the BLEU scores of Qwen-72B-chat by 122% on zh2za and by 67% on za2zh over the original DiPMT. We will further discuss how each design in DiPMT++ contributes to the overall performance in Section 6.
@@
Model Scale. In terms of model scales, we observe that the performance is steadily improved with the increase of model parameters for LLaMA-2 and Qwen. Since Qwen has a better Chinese capability than the English-centric LLaMA-2, a 14B Qwen model can outperform a 70B LLaMA-2. GPT-4 outperforms all the other models, demonstrating its excellent reasoning ability. It is worth noting that the open-source Qwen-72B-chat performs comparably to the closed-source GPT-4 on the zh2za task, which is an encouraging result for more transparent and reproducible research on low-resource NLP.
5.3 Additional Experiments on Kalamang

### v2

@@ Abstract
Existing large language models struggle to support numerous low-resource languages, particularly the extremely low-resource ones, for which there is minimal training data available for effective parameter updating. We thus investigate whether LLMs can learn a new language on the fly solely through prompting. To study this question, we collect a research suite for Zhuang, a language supported by no LLMs currently. We introduce DiPMT++, a framework for adapting LLMs to unseen languages by in-context learning. Using a dictionary and 5K parallel sentences only, DiPMT++ significantly enhances the performance of GPT-4 from 0 to 16 BLEU for Chinese-to-Zhuang translation and achieves 32 BLEU for Zhuang-to-Chinese translation. We also validate the effectiveness of our framework on Kalamang, another unseen language. Furthermore, we demonstrate the practical utility of DiPMT++ in aiding humans in translating completely unseen languages, which could contribute to the preservation of linguistic diversity.
@@ 1 Introduction
Existing large language models (LLMs) provide robust support for many high-resource languages, but their support for numerous low-resource languages is limited Ahuja et al. (2023). To adapt LLMs to low-resource languages, continual pre-training or adaptors are commonly employed Pfeiffer et al. (2020); Yong et al. (2023). However, a corpus of only a few thousand sentences is insufficient to update the model parameters effectively for learning extremely low-resource languages Joshi et al. (2020). Considering the inductive and mimicking capabilities of LLMs, an interesting research question arises: Can LLMs learn a new low-resource language on the fly solely through prompting? This learning paradigm could enable more efficient utilization of limited resources and holds significant potential in the preservation and education of underrepresented languages.
@@
To explore this question, we choose Zhuang (ISO 639-1: za; ISO 639-3: zha), an extremely low-resource language, as our focus. It is the language spoken by the Zhuang people of Southern China. [^1] See more information about Zhuang in Appendix A. There are no open-source natural language processing (NLP) datasets in Zhuang, and existing LLMs do not support this language. Therefore, we curate ZhuangBench, a research suite for Zhuang, comprising a dictionary, a parallel corpus of 5K Zhuang-Chinese sentences, and a machine translation test set. ZhuangBench is not only a valuable linguistic resource for this extremely low-resource language but also a challenging benchmark for LLMs, with which we can investigate how models learn an entirely new language.
We focus on the machine translation (MT) task in ZhuangBench. Previous research such as DiPMT Ghazvininejad et al. (2023) has explored translation in low-resource languages via prompting, providing translations for rare words with existing dictionaries. This method is tailored for languages where LLMs already possess fundamental capabilities, especially basic knowledge for syntax, which might be learned in the form of $n$-gram language modeling. However, when confronted with an entirely new language where LLMs need to acquire its vocabulary and grammar from scratch, previous prompting methods hardly work.
In this work, we introduce DiPMT++, a framework to efficiently adapt LLMs to an unseen language via in-context learning (ICL). Built upon DiPMT, our method provides models with the meanings of words appearing in the source sentence, as illustrated in Figure 1. Given the inherent incompleteness of the dictionary for low-resource languages, we enhance the lexical coverage by revisiting traditional techniques like bilingual lexicon induction and synonym expansion. To aid models in grasping basic syntax, we retrieve closely related exemplars from parallel corpora to construct the ICL prompt.
@@
To investigate the applicability of DiPMT++ in realistic scenarios, we conduct a user study of unseen language translation. We recruit participants who have no knowledge of Zhuang, and ask them to conduct translation with the given linguistic sources. Experiments reveal that if we assist humans with DiPMT++, their translation improves in both quality and efficiency. This shows our framework’s great potential in preserving endangered languages.
Our contributions are as follows:
$\bullet$ We present ZhuangBench, a challenging benchmark for LLMs to translate an unseen language with limited linguistic resources.
$\bullet$ We develop an ICL framework DiPMT++ for on-the-fly language learning, which has proven effective on two benchmarks, and explore more strategies for enhancing LLMs’ acquisition of lexical and syntactic knowledge.
$\bullet$ We showcase that DiPMT++ can assist humans in translating unseen languages, which could benefit the preservation of linguistic diversity.
Our code and data are available to the public [^2] https://github.com/luciusssss/ZhuangBench.
@@ 2 Related Works
Adapting LLMs to Low-Resource Languages
Continual pretraining on monolingual texts is a common practice to adapt LLMs to low-resource languages Yong et al. (2023); Zhang et al. (2023b). Techniques such as MAD-X Pfeiffer et al. (2020, 2021) and LoRA Hu et al. (2021) are used to improve the training efficiency. Yang et al. (2023) design a high- to low-resource curriculum for multilingual continual pertaining. Purkayastha et al. (2023) attempt to romanize the unseen scripts for language adaptation.
Another line of work directly adapts instruction-tuned LLMs through supervised fine-tuning (SFT) by constructing cross-lingual instructions Cahyawijaya et al. (2023) or leveraging pivot languages Zhang et al. (2023c). Yong et al. (2023) find that SFT is sometimes more efficient for language adaptation than continual pretraining.
Unlike previous works, we explore the possibility of on-the-fly language adaptation through prompting given minimal linguistic resources.
Machine Translation with LLMs
Recent works demonstrate the effectiveness of prompting an LLM with a few MT examples Zhang et al. (2023a); Vilar et al. (2023); Garcia et al. (2023). Following them, researchers explore different strategies for ICL exemplar selection, from the perspectives of term recall Agrawal et al. (2023), knowledge relevance He et al. (2023), and cultural awareness Yao et al. (2023). However, these strategies are primarily tailored for high-resource languages.
Robinson et al. (2023) show that LLMs still fail to translate low-resource languages properly. Existing methods designed for low-resource languages include incorporating dictionaries Ghazvininejad et al. (2023); Elsner and Needle (2023), using high-resource languages as pivots Jiao et al. (2023), and refining results from smaller MT systems Cheng et al. (2023). In contrast to them, our work tackles an even more challenging scenario: translating languages entirely unseen by LLMs.
@@ 3 Dataset: ZhuangBench
Dictionary
The Zhuang-Chinese dictionary is collected from an online dictionary site [^3] https://zha_zho.en-academic.com/, with 16,031 Zhuang words. The average number of senses for each word is 1.4. We also convert it to a Chinese-Zhuang dictionary with 13,618 Chinese words, with an average of 2.2 translations per word.
Parallel Corpus
The parallel corpus contains 4,944 Zhuang-Chinese sentence pairs from multiple sources. 2,135 pairs are obtained from the Chinese and Zhuang versions of the Government Work Reports in China, where we map each Chinese sentence to its corresponding Zhuang translation. 2,127 pairs are collected from Zhuang textbooks. The remaining 682 pairs are example sentences in the dictionary. See more statistics in Appendix B.
Translation Test Set
We provide a hold-out test set with 200 sentence pairs for evaluation. It includes instances of three difficulty levels: 75 easy instances, 60 medium ones, and 65 hard ones.
The easy subsets are composed of sentences from an elementary textbook. Other subsets are collected from the official Zhuang Language Proficiency Test (Vahcuengh Sawcuengh Suijbingz Gaujsi, V.S.S.G.) in China, which is divided into three levels: elementary, intermediate, and advanced. The translation instances from the elementary and intermediate levels of V.S.S.G. form our medium subset while those from the advanced level constitute our hard subset. In Appendix B, we list statistics and examples for each level. The three subsets of different difficulties exhibit a gradual increase in vocabulary coverage and sentence complexity.
@@ 4 Method
We introduce DiPMT++, a language-agnostic framework to adapt LLMs to an unseen language efficiently. It can serve as a strong baseline for the machine translation task in ZhuangBench.
@@ 4.1 Preliminary: DiPMT
DiPMT Ghazvininejad et al. (2023) is a prompting-based method for low-resource language translation. Given a source sentence, the model looks up in the dictionary for the meaning of rare words and adds them to the prompt directly with the format in this context, the word “[source word]” means “[target word]”. The prompt is adopted in an in-context learning manner, with $k$ demonstrations before the current testing instance.
DiPMT is designed for languages that current models perform moderately well (10 - 30 BLEU scores). The authors claim that DiPMT is not suitable for languages where current models perform poorly (< 10 BLEU points), as they assume that the performance for those is too low to expect reasonable translations even when incorporating external information.
@@ 4.2 Our Method: DiPMT++
One may ask what we can do to help LLMs understand those extremely low-resource languages, for which we only have a parallel corpus of a few thousand sentences. Contrary to the pessimistic view of Ghazvininejad et al. (2023), we hypothesize that it is possible to teach LLMs a new language solely through prompting, considering LLMs’ impressive ability to infer and mimic. We make extensions to the DiPMT framework so that it can be applied to extremely low-resource languages. Leveraging the powerful reasoning capabilities of LLMs, we propose DiPMT++, a new method that allows LLMs to understand a completely new language with minimal resources.
Following DiPMT, we cast the on-the-fly machine translation as an ICL task and incorporate knowledge from bilingual dictionaries. DiPMT++ makes two key modifications to DiPMT as follows. See the prompt template in Appendix C.
Improved Lexical Coverage
DiPMT only provides meanings for less frequent words in the sentence. For an unseen language, we need to provide translations for as many words in the sentence as possible to the model. However, not all words can be found in the dictionary for a low-resource language. DiPMT++ attempts to improve lexical coverage of the prompt by revisiting traditional statistical methods and linguistic resources. Specifically, we use the following three strategies:
@@
$\bullet$ Bilingual Lexicon Induction: Even a small-scale parallel corpus might contain words not included in the dictionary. Traditional statistical methods such as GIZA++ Och and Ney (2003) can efficiently mine bilingual lexicon from the corpus, which could complement the dictionary.
@@
Syntactically-Informed Exemplar
In DiPMT, the exemplars are fixed and have limited relevance to the testing instance. Their function is only to demonstrate the task. It hardly works when the LLMs have little knowledge about the grammar of an unseen language. DiPMT++ attempts to dynamically select exemplars with higher relevance and encourage models to infer elementary syntactic information from the exemplars. For a testing instance, its exemplars are retrieved from a parallel corpus. Although there are advanced retrievers such as DPR Karpukhin et al. (2020), they require additional training and do not support extremely low-resource languages like Zhuang. Therefore, we apply BM25 Robertson et al. (2009), a language-agnostic retrieval algorithm, for exemplar retrieval.
Model Chinese $\to$ Zhuang (BLEU / chrF) Zhuang $\to$ Chinese (BLEU / chrF)
@@
Table 1: Performance of different methods on the test set of ZhuangBench. We use 3-shot exemplars for prompting-based methods. The best scores are made bold, with the second underlined.
@@ 5.1 Experimental Setup
Backbone Models
We use three types of models as the backbone of DiPMT++: (1) Llama-2-chat Touvron et al. (2023), an open-source English-centric model, (2) Qwen-chat Bai et al. (2023), a bilingual model for English and Chinese, and (3) GPT-3.5 and GPT-4 OpenAI (2023), two commercial multilingual models [^4] The versions of the OpenAI APIs are gpt-3.5-turbo-0125 and gpt-4-0125-preview..
Baselines
We adopt a variety of baselines for comparison. (1) Finetune: Finetuning models with parallel sentences. (2) Direct: Directly asking LLMs to perform translations without providing ICL exemplars, which, to some extent, reflects whether the LLM already knows the language. (3) DiPMT Ghazvininejad et al. (2023): Using the original design of DiPMT for LLM prompting.
Metrics
We use BLEU Papineni et al. (2002) and chrF Popović (2015), implemented by sacreBLEU Post (2018). BLEU is a word-level metric while chrF focuses on the character level.
Tokenization
All the models used in our experiments support Chinese. Because Zhuang adopts a Latin script, these models can tokenize Zhuang texts into subwords, or characters at least, without producing UNK. For example, Llama-2’s tokenizer tokenizes Liuz coengmingz. into [‘_Li’, ‘uz’, ‘_co’, ‘eng’, ‘ming’, ‘z’, ‘.’].
@@ 5.2 Results and Analyses
In Table 1, we report the results on the Chinese-to-Zhuang (zh2za) and Zhuang-to-Chinese (za2zh) translation task of ZhuangBench. See samples of output from different models in Appendix D.
Finetuning vs. Prompting
@@
Finetuning pretrained models is a common practice for low-resource machine translations Adelani et al. (2022). Fintuned smaller models like mT5 still have BLEU scores close to zero for Zhuang [^5] Although mT5-large achieves 15.1 chrF on zh2za, the model outputs are almost non-sense, as shown by the example in Appendix D. As Zhuang uses a Latin script with 26 characters, even a meaningless sequence would likely have a non-zero chrF score.. Through finetuning, Llama-2-7B-chat can develop a basic understanding of Zhuang. It performs particularly well on the easy subset of zh2za translation but still struggles with more challenging instances. We refrain from finetuning larger models due to the substantial computational resources required.
@@
Compared to the high expense of finetuning, prompting with DiPMT++ requires no training while delivering comparable or even superior performance when combined with larger models. Prompting Llama-2-7B-chat with DiPMT++ only lags finetuning by 0.6 BLEU on zh2za and by 1.6 BLEU on za2zh. Despite having little prior knowledge about Zhuang, as evidenced by their poor performance in direct prompting, GPT-3.5 and GPT-4 achieve excellent results on the translation tasks of varying difficulty levels with the assistance of DiPMT++. For instance, GPT-4 paired with DiPMT++ achieves a remarkable BLEU score of 31.9 on za2zh, which might be qualified for practical use.
DiPMT vs. DiPMT++
@@
We compare DiPMT and DiPMT++ with Qwen-chat as the backbone model. Although useful for mid-source languages, the original DiPMT has limited ability to assist LLMs in understanding a completely new language. Even Qwen-72B-chat, the largest version of Qwen-chat, achieves only 5.1 BLEU on Chinese-to-Zhuang translation. After introducing two simple extensions, DiPMT++ activate the reasoning ability of LLMs and greatly boost the performance. For example, DiPMT++ increases the BLEU scores of Qwen-72B-chat by 122% on zh2za and by 67% on za2zh over the original DiPMT. We will further discuss how each design in DiPMT++ contributes to the overall performance in Section 6.
Model Scale
@@
Regarding model scales, we observe that the performance steadily improves with the increase of model parameters for Llama-2 and Qwen. Since Qwen has a better Chinese capability than the English-centric Llama-2, a 14B Qwen model can outperform a 70B Llama-2. GPT-4 outperforms all other models, demonstrating its excellent reasoning ability. It is worth noting that the open-source Qwen-72B-chat performs comparably to the closed-source GPT-4 on the zh2za task, which is an encouraging result for more transparent and reproducible research on low-resource NLP.
5.3 Additional Experiments on Other Languages
Another Unseen Language: Kalamang